OpenStudy (anonymous):

What is the solution to the equation 4x + 2(x − 3) = 3x + x − 12?

OpenStudy (pinklion23):

what do you think the answer is

OpenStudy (pinklion23):

x = -3 :/ can i have medal too XD

OpenStudy (hungergames13):

-24

OpenStudy (hungergames13):

-24 is the solution to the equation. :)

OpenStudy (gabbyalicorn):

4x+2(x−3)=3x+x−12 Simplify both sides of the equation. 4x+2(x−3)=3x+x−12 Simplify 6x−6=4x−12 Subtract 4x from both sides. 6x−6−4x=4x−12−4x 2x−6=−12 Add 6 to both sides. 2x−6+6=−12+6 2x=−6 Divide both sides by 2. 2x2=−62 x=−3 Answer: x=−3